# Sparse Repellency for Shielded Generation in Text-to-Image Diffusion Models

### v2

@@ 4.2 Benchmark
We first examine whether adding SPELL post-hoc increases the diversity of trained diffusion models. To this end, we quantitatively compare each diffusion model to the same model run with the same random generation seeds but with SPELL. In particular, we use intra-batch repellency together with repellency from previously generated batches, to enable repellency across the up to 128 images per prompt/class. Table 1 shows that SPELL consistently increases the diversity, both in terms of recall and Vendi score, across all text-to-image and class-to-image diffusion models. This demonstrates that SPELL works independently from the model architecture and the space the models diffuse in (RGB space for Simple Diffusion, VAE space for all others). Coverage remains within -1% to +2% of its original value in all models except SD3. The difference between coverage and recall is that coverage uses a more tight neigborhood radius to determine whether an image of the original dataset is covered by the generated ones. In other words, coverage measures a form of dataset match, which can counter-intuitively be decreased by more diverse outputs if the diversity takes different forms or is higher than in the reference dataset. This stands out most for SD3, which was not trained on the reference datasets ImageNet-1k/CC12M. We find that SPELL correctly helps SD3 generate images that are generally more diverse, as evidenced by the 26% and 37% increases in the reference-dataset-free Vendi score, but in other attributes than in the reference datasets, explaining the decrease in coverage. Precision and density decrease slightly in five out of the six models, but stay within -3% to +2% of the original values, the exception again being SD3. This tradeoff between diversity and precision is common in the literature (Kynkäänniemi et al., 2024; Sadat et al., 2024; Corso et al., 2024), and we show in the next section that our method provides more favorable Pareto fronts than alternative recent methods. This tradeoff improves the overall $\text{FD}_{\text{DINOv2}}$ score considerably in Latent Diffusion and MDTv2, while staying within 3% of the original value on Simple Diffusion and EDMv2, and increasing on SD3. Overall, we find that SPELL increases the diversity considerably across all models, with only minor tradeoffs in precision.
@@
Figure 4: Latent Diffusion on CC12M. The hyperparameters of diversity methods allow trading off image quality (x-axes) for diversity metrics (y-axes). Our repellency provides a better trade-off than other concurrent approaches. IG is not visible on all plots as it strongly decreases image quality.
@@ 4.3 Comparison to Other Diversity-inducing Methods
We now take a closer look at SPELL’s hyperparameter, the repellency radius $r$. Figure 3 shows that we can use repellency in two ways. A low radius ($r=15$) increases the diversity (Vendi score and recall) without compromising on precision. Intuitively, such a small radius only serves to prevent generating the nearly-same image twice. The resulting output set also better reflects the true image distribution, yielding an enhanced $\text{FD}_{\text{DINOv2}}$. The second option is to further increase the repellency radius, which as Figure 3 shows further enhances diversity at the cost of precision.
There are other current proposals in the literature that enable to control this trade-off. Namely, Interval Guidance (Kynkäänniemi et al., 2024, IG) only applies CFG in a limited time interval in the middle of the backward diffusion. Condition-annealed diffusion sampling (Sadat et al., 2024, CADS) noises the text or class condition that guides the CFG, lowering the noise in later timesteps. Particle Guidance (Corso et al., 2024, PG) adds a gradient potential to the backward diffusion such that the diversity of a batch as a whole is increased. We reimplement and tune the hyperparameters of all of these approaches and compare them against SPELL.
Figure 4 shows that SPELL achieves a more favorable trade-off curve than each of these methods, in three different diversity vs quality Pareto fronts, namely recall vs precision, coverage vs density, and Vendi score vs CLIP Score. One reason for this is SPELL’s sparsity. SPELL only intervenes if images are going to be too similar to one another. Due to its ReLU weighting, it remains exactly zero if an image’s trajectory is already headed to a diverse output. This leads to increased diversity while leaving high-precision images unchanged. We study this sparsity further in the following section.
@@ 4.4 Sparsity Analysis
Figure 5: (a) The gradient that SPELL adds is only a fraction of the magnitude of the diffusion’s score, thus adjusting it without drowning it out. (b) Repellency happens primarily in early backwards steps ($t\in[0.6,1.0])$ and then often remains zero, thus making it sparse. Overcompensation allows finishing the repellency earlier, whereas runs with high repellency radius repel longer. Latent Diffusion trajectories with repellency from both previous generations and intra-batch.
We focus in this section on the dynamics of SPELL interventions and investigate when and how SPELL corrective terms arise. Figure 5 tracks the magnitude of the SPELL correction vector $\|\Delta\|_{2}$ normalized by that of the diffusion score vector $\|\nabla\log p_{t}(x_{t}|y)\|_{2}$. We track this relative magnitude throughout 452 backward trajectories for 50 prompts of CC12M. Because our experiment considers both intra-batch repellency and repellency, as in Table 1, the SPELL correction is equal to the addition of past comparison terms (Equation 5) and within batch corrections (Equation 8). Appendix F gives corresponding plots for further setups. First, we find that repellency only adds a small correction term in most cases. Its magnitude is most often less than 5% of the magnitude of the diffusion score and never exceeds 35%. This explains why our repellency does not reduce image quality or introduce artifacts. A second reason for this is that SPELL corrections happen mostly in the early stages of the backward diffusion, which literature claims to be when the rough image is outlined, rather than in late steps, where the image is refined (Biroli et al., 2024; Kynkäänniemi et al., 2024). Recall that the backwards diffusion starts at $t=1$ and outputs the final image at $t=0$. Figure 5 shows that at $t=0.8$, only 40% of the trajectories have a non-zero repellency term anymore, further declining to 21% at $t=0.6$. If we impose a higher repellency radius, the repellency acts for longer. Especially in this cases, adding overcompensation helps. As intended, the repellency strength is increased and in return stops earlier. These stops are often final: The repellency stays zero for the remainder of the generation, verifying that the trajectories do not bounce back into the repellency radii, as shown in Appendix F, and reaffirming SPELL’s sparsity.
4.5 Qualitative Example of SPELL Interventions
Figure 6 shows 16 images generated iteratively using SD3 with and without SPELL interventions. Images are generated one by one $(B=1)$, and when generating the $i+1$-th image, SPELL repels from the reference set of all images $1$ to $i$ it has generated thus-far. Images are highlighted in orange if SPELL enforced changes during their generation trajectory. When SPELL does not intervene, the SD3 + SPELL image (bottom row) coincides with the SD3 output (top row), since we use the same seeds. For images 1 to 3, SPELL did not intervene as it did not detect similarities when generating the 2nd image when compared to the 1st, nor when generating the 3rd w.r.t. 1st and 2nd. On the contrary, the 4th image was expected to come out too close to the 3rd at some time during generation, triggering SPELL interventions that altered ultimately the ball color, dog, and background. As more images are added to the reference set, SPELL intervenes more often. For example, image 14 is headed towards an image of a dog on a grassy ground with trees in the background, which is too similar to multiple previous images. Hence, SPELL intervenes and leads the diffusion trajectory to an entirely new mode, with both a new dog race and a previously unseen tartan surface. A similarly strong intervention happens in image 15. Still, SPELL’s sparsity means that it does not always intervene, even when there are already many shielded images. For example, image 16 is novel enough to remain unchanged.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption{3}{4}{3, 6}{4}{2, 8}{4, 8}{4, 6, 7, 8}{2, 6}{2, 3, 4, 5, 6, 12}  {2, 4, 6, 8}1122334455667788991010111112121313141415151616
@@ 4.5 Qualitative Example of SPELL Interventions
Figure 6: Images for the prompt “a dog plays with a ball” generated one-by-one ($B=1$) with SD3 (top) and SD3 + SPELL (bottom), using the same random seeds. SPELL intervenes on 10 of the 16 images, marked with orange borders, because they are too similar to previously generated images. Images that SPELL repelled from are listed below each image. Early images are adjusted less often and mostly in details because they are still novel enough. Later images repel from more previous images and more strongly to ensure they are different enough.
4.6 Image Protection Benchmark
We scale SPELL to a repellency set of 1.2 million ImageNet-1k train images via approximate nearest neighbor search (Douze et al., 2024), with the goal of generating novel images. We generate 50k images, track how often they fall into a protected shield, as well as the precision, recall, and runtime. This is similar to machine unlearning for generative models (Wang et al., 2024a; Liu et al., 2024), with the main differences that we wish to shield a set of specific images rather than global concepts (Wu et al., 2024; Park et al., 2024), and that SPELL is a training-free intervention.
Table 2 shows that 7.6% of the 50k images that MDTv2 generates without SPELL are within an $L_{2}$ distance of $r=60$ of their nearest neighbor on ImageNet. Figure 7 shows examples and verifies that such images indeed are nearly copies of existing images. Adding SPELL with $r=60$ reduces this rate down to 0.16%. Figure 7 shows that the images are indeed changed in such a way that they are not too close to their ImageNet neighbors anymore. Searching for shields in less Voronoi cells of the approximate nearest neighbor algorithm allows to speed up the generation time, at the cost of a catching less shields. This runtime, and the fact that there are still images within the shields, can largely be attributed to the approximate next neighbor search algorithm. Further improvements in $L_{2}$ based neighbor search techniques will further increase SPELL’s protection rate and compute overhead. Expectedly, the recall decreases when repelling from all training images, because validation images may fall into the shield radius around train images. However, the precision remains largely unaffected. This demonstrates again that SPELL does not introduce visual artifacts, even when repelling from many images. To add qualitative evidence to this point, we show a random set of images where SPELL intervened in Appendix H. Finally, the last two images in Figure 7 give more insight into the workings of the $L_{2}$ similarity in the VAE latent space that MDTv2 diffuses in. Apparently, image distances inside the VAE space encode a visual similarity where images with similar colors and compositions are close to one another. We leave applications of SPELL with more semantic similarities, e.g., by comparing not the expected image outputs but the DINOv2 embeddings of expected image outputs, for future works.
Table 2: SPELL almost always generates images outside the shields of the ImageNet-1k train set. Searching for shields in more Voronoi cells during the nearest neighbor lookup improves the protection rate at a higher inference-time runtime. The runtime is reported on a single A100-40GB GPU.
@@ 4.6 Image Protection Benchmark
Model Searched cells Generated images too close to ImageNet neighbors $\downarrow$ Precision $\uparrow$ Recall $\uparrow$ Time per image (s) $\downarrow$
EDMv2 7.60% 0.792 0.242 2.434
+ SPELL 1 1.08% 0.792 0.181 4.633
+ SPELL 2 0.55% 0.788 0.175 6.057
+ SPELL 3 0.33% 0.777 0.162 7.790
+ SPELL 5 0.22% 0.771 0.163 9.949
+ SPELL 10 0.16% 0.768 0.160 13.545
EDMv2samplesRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionImageNetneighborRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionEDMv2 + SPELLRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionImageNetneighborRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption

### v3

@@ 4 Related Works
Chen et al. (2024) is also closely related. Here, the authors apply anti-memorization guidance to pretrained models through three terms: desspecification; caption de-duplication and dissimilarity guidance. Similar to Sadat et al. (2024), desspecification adjusts the CFG scale dependent on nearest neighbor in the training data to $\mathbb{E}[\mathbf{X}_{0}|x_{t}]$. Caption de-duplication adds a negative guidance term (Liu et al., 2022) based on the network applied to the caption of the nearest neighbor. Finally, dissimilarity guidance applies an additional guidance term based on similarity between $\mathbb{E}[\mathbf{X}_{0}|x_{t}]$ and nearest neighbor search. Our approach is most similar to dissimilarity guidance using multiple nearest neighbors and carefully chosen guidance scale to encourage generated samples to be outside a radius of reference points.
Since initial release of our work, Koulischer et al. (2024) has provided more theoretical grounding for weighting negative guidance (Liu et al., 2022) based on an estimated class-based likelihood. Kim et al. (2025) demonstrates negative guidance using the empirical score of a reference set of images. Unlike aforementioned guidance based approaches, Thornton et al. (2025) introduces a gradient-free SMC-based sampler that encourages diversity by negatively weighting the energy in a Feynman Kac potential, demonstrated only on toy examples.
5 Experiments
We now show that SPELL increases the diversity of modern text-to-image and class-conditional diffusion models (Section 5.2), with a better trade-off than other recent diversity methods (Section 5.3). We quantify the sparsity of SPELL interventions in Section 5.4. In Section 5.6, we demonstrate SPELL’s scalability and a new use-case, shielded generation, by generating novel ImageNet images while shielding all 1.2 million ImageNet-1k train images.
5.1 Experimental Setup
In the class-to-image setup, we use Masked Diffusion Transformers (MDTv2) (Gao et al., 2023), EDMv2 (Karras et al., 2024), and Stable Diffusion 3 Medium (SD3) (Esser et al., 2024), three recent state-of-the-art diffusion models. We use the pretrained model checkpoints to generate 50,000 256x256 images of ImageNet-1k classes(Deng et al., 2009) without and with SPELL and compare them to the original ImageNet-1k images. We use the validation dataset as a comparison, since we will conduct experiments that repel from the training dataset in Section 5.6, which would render comparisons to the training dataset meaningless. Given the text-to-image setting, we follow (Esser et al., 2024) using “a photo of a class_name” as caption.
In our text-to-image setup, we use SD3, Latent Diffusion (Rombach et al., 2022), and RGB-space Simple Diffusion (Hoogeboom et al., 2023) in resolution 256x256. For the latter two, we use the checkpoints of Gu et al. (2023). Details on hyperparameters are provided in Appendix D. We evaluate these models on CC12M (Changpinyo et al., 2021), a dataset of (caption, image) pairs, with captions ranging between 15 and 491 characters. As we aim to investigate whether diffusion models with and without SPELL capture the full diversity of images related to each prompt, we create a sub-dataset of captions with multiple corresponding images, which gives a ground-truth target diversity. This gives a one-to-many setup with 5000 captions and 4 to 128 images each (in total 41,596 images). We explain the construction of this dataset in Appendix E.
To evaluate diversity, we track the recall (Kynkäänniemi et al., 2019), coverage (Naeem et al., 2020), and Vendi score (Friedman & Dieng, 2023). To evaluate image quality, we use precision (Kynkäänniemi et al., 2019) and density (Naeem et al., 2020). We track these metrics per class/prompt and average across classes/prompts. To measure whether the generated images match the true image distributions, we use the marginal FID (Heusel et al., 2017) and the marginal Fréchet Distance with DINOv2 features ($\text{FD}_{\text{DINOv2}}$, Stein et al. (2024); Oquab et al. (2024)).
@@ 5.1 Experimental Setup
Figure 4: Latent Diffusion on CC12M. The three plots on the left highlight how the hyperparameters of diversity methods trade off image quality (x-axes) and diversity metrics (y-axes). SPELL provides a better trade-off than other concurrent approaches. In the rightmost plost, highlighting 2 quality metrics, SPELL also shines. IG is not visible on all plots as it strongly decreases image quality.
5.2 Benchmark
We first examine whether adding SPELL post-hoc increases the diversity of trained diffusion models. To this end, we compare each diffusion model to the same model run with the same random generation seeds but with SPELL. In particular, we use intra-batch repellency together with repellency from previously generated batches, to enable repellency across the up to 128 images per prompt/class.
Table 1 shows that SPELL consistently increases the diversity, both in terms of recall and Vendi score, across all text-to-image and class-to-image diffusion models. This demonstrates that SPELL works independent of the model architecture and the space the models diffuse in (RGB space for Simple Diffusion, VAE space for all others). Coverage remains within -1% to +2% of its original value in all models except SD3. The difference between coverage and recall is that coverage uses a more tight neigborhood radius to determine whether an image of the original dataset is covered by the generated ones. In other words, coverage measures a form of dataset match, which can counter-intuitively be decreased by more diverse outputs if the diversity takes different forms or is higher than in the reference dataset. This stands out most for SD3, which was not trained on the reference datasets ImageNet-1k/CC12M. We find that SPELL correctly helps SD3 generate images that are generally more diverse, as evidenced by the 26% and 37% increases in the reference-dataset-free Vendi score, but in other attributes than in the reference datasets, explaining the decrease in coverage. Out of the six experiments, precision and density decrease very slightly in three of them, increase for one, and decrease more clearly in 2 (when using SD3). This tradeoff between diversity and precision is common in the literature (Kynkäänniemi et al., 2024; Sadat et al., 2024; Corso et al., 2024), and we show in the next section that SPELL provides more favorable Pareto fronts than alternative recent methods. This tradeoff improves the overall $\text{FD}_{\text{DINOv2}}$ score considerably in Latent Diffusion and MDTv2, while staying within 3% of the original value on Simple Diffusion and EDMv2, and increasing on SD3. Overall, we find that SPELL increases the diversity considerably across all models, with only minor tradeoffs in precision.
@@ 5.3 Comparison to Other Diversity-inducing Methods
We now consider SPELL’s hyperparameter, the repellency radius $r$. Intuitively, a small radius will only prevent generating a nearly-same image twice, which increases the diversity without compromising on the closeness to the train dataset (precision/density). Choosing to further increase the repellency radius $r$ should add more diversity while trading-off precision, as SPELL pushes the repellency trajectories to explore novel modes further outside the train distribution. We find in Figure 26 of Appendix M that this intended effect plays out in practice. We also find that there is a sweet-spot for $r$ where the diversified samples better reflect the true image distribution than the base model, enhancing $\text{FD}_{\text{DINOv2}}$ and diversity without decreasing precision.
There are some recent methods enable controlling similar diversity-precision trade-offs. Namely, Interval Guidance (Kynkäänniemi et al., 2024, IG) applies CFG in a limited time interval in the middle of the backward diffusion. Condition-annealed diffusion sampling (Sadat et al., 2024, CADS) noises the text or class condition that guides the CFG, lowering the noise in later timesteps. Closer in spirit to our proposal, particle guidance (Corso et al., 2024) adds a gradient potential to the backward diffusion at every timestep, such that the intra-batch diversity is increased. We reimplement and tune these baselines (Appendix F) to compare them against SPELL.
Figure 4 shows that SPELL achieves more favorable trade-off curves in three different diversity vs quality Pareto fronts (recall vs precision, coverage vs density, and Vendi score vs CLIP Score) as well as in a prompt-adherence vs quality Pareto front (CLIP vs $\text{FD}_{\text{DINOv2}}$). At low repellency radii $r$, the diversity is achieved without reducing the CLIP score. In Appendix K, we find that SPELL generates diverse images both for short and for longer, more specific prompts.
One reason for SPELL’s improved performance is sparsity. While other methods change the diffusion trajectories at every timestep and every image, due to it’s ReLU weighting, guidance is not applied if the diffusion trajectories are already heading to a diverse outputs. This leads to increased diversity while leaving high-precision images unchanged. We study this sparsity further in the following section.
@@ 5.4 Sparsity Analysis
Figure 5: (a) The gradient that SPELL adds is only a fraction of the magnitude of the diffusion’s score, thus adjusting it without drowning it out. (b) Repellency happens primarily in early backwards steps ($t\in[0.6,1.0])$ and then remains zero, thus making it sparse. Overcompensation allows finishing the repellency earlier, whereas runs with high repellency radius repel longer. Latent Diffusion trajectories with intra-batch and previous-batch repellency.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption{2}{3}{2, 5}{3}{1, 7}{3, 7}{3, 5, 6, 7}{1, 5}{1, 2, 3, 4, 5, 11}  {1, 3, 5, 7}11223344556677889910101111121213131414
@@
Figure 6: High-resolution images for the prompt “a dog plays with a ball” generated one-by-one ($B=1$) with SD3 (top) and SD3 + SPELL (bottom), using the same random seeds. While SPELL does not intervene on the first two images, as they are different enough, it intervenes on the 10 marked with orange borders, as they are too similar to previously generated images. This changes the output from what it would have been without SPELL (at the top) to slightly or completely novel images at the bottom.
This section investigates the dynamics of when and how strongly SPELL’s corrective interventions arise. Figure 5 tracks the magnitude of the SPELL correction vector $\|\Delta\|_{2}$ normalized by that of the diffusion score vector $\|\nabla\log p_{t}(x_{t}|y)\|_{2}$. We track this relative magnitude throughout 452 backward trajectories for 50 prompts of CC12M with both intra- and inter-batch repellency (Equations 5 and 7). Appendix H adds further setups with similar results. We find that the repellency correction is typically small. Its magnitude is most often less than 5% of that of the diffusion score and never exceeds 35%. This explains why our repellency does not reduce image quality or introduce artifacts. A second reason for this is that SPELL corrections happen mostly in the early stages of generation, which literature claims to be when the rough image is outlined, rather than in late steps, where the image is refined (Biroli et al., 2024; Kynkäänniemi et al., 2024). Recall that the backwards diffusion starts at $t=1$ and outputs the final image at $t=0$. Figure 5 shows that at $t=0.8$, only 40% of the trajectories have a non-zero repellency term, further declining to 21% at $t=0.6$. If we impose a higher repellency radius $r$, the repellency acts for longer. Especially in this case, adding overcompensation helps. As intended, the repellency strength is increased and in return stops earlier. These stops are often final: The repellency stays zero for the remainder of the generation, verifying that the trajectories do not bounce back into the repellency radii, as shown in Appendix H, and reaffirming SPELL’s sparsity.
5.5 Qualitative Example of SPELL Interventions
This section shows what SPELL’s correction terms change in practice. Figure 6 shows 14 images in 1024x1024 resolution, generated iteratively using SD3 with and without SPELL interventions. Images are generated one by one $(B=1)$, and when generating the $i+1$-th image, SPELL repels from the reference set of all images $1$ to $i$ it has generated thus-far. Images are highlighted in orange if SPELL enforced changes during their generation trajectory. When SPELL does not intervene, the SD3 + SPELL image (bottom row) coincides with the SD3 output (top row), since we use the same seeds. For images 1 and 2, SPELL did not intervene as it detected that the 2nd trajectory was heading to an image that was novel enough from the 1st. The 3rd image was expected to come out too close to the 2nd at some time during generation, triggering SPELL to alter the background, ball color, and details on the dog. As more images are added to the reference set, SPELL intervenes more often. For example, image 13 is guided away from an image of a dog on a grassy ground with trees in the background, which images 2, 3, 5, and 11 already show, and explores an entirely novel mode, with both a new dog race and a previously unseen surface. A similarly strong intervention happens in image 14. SPELL’s sparsity means that it it not always applied, even when there are already many shielded images. This is the case for image 8, which is novel enough to remain unchanged. Note that none of the images has visual artifacts due to SPELL’s early and sparse interventions. Appendices J and N present hundreds of images generated with SPELL on further models, affirming this finding.
Table 2: Without SPELL, 7.60% of EDMv2’s generated images are too close to the protected ImageNet-1k train set. Adding SPELL reduces this rate, and the more inference-runtime is spent on searching Voronoi cells (1, 2, 3, 5, or 10), the better the protection becomes. The runtime is reported on a single A100-40GB GPU.
@@ 5.5 Qualitative Example of SPELL Interventions
Model Gen. images within shields $\downarrow$ Prec. $\uparrow$ Recall $\uparrow$ Time per image (s) $\downarrow$
EDMv2 7.60% 0.792 0.242 2.43
+ SPELL-1 1.08% 0.792 0.181 4.63
+ SPELL-2 0.55% 0.788 0.175 6.06
+ SPELL-3 0.33% 0.777 0.162 7.79
+ SPELL-5 0.22% 0.771 0.163 9.94
+ SPELL-10 0.16% 0.768 0.160 13.54
5.6 Image Protection at Scale
Last, we present a second use-case of shielded generation, where the goal is to create images that are sufficiently novel from a given large set of images. In particular, we scale SPELL to shield all 1.2 million ImageNet-1k train images by employing approximate nearest neighbor search (Douze et al., 2024, see Appendix I). We then generate 50k images and track how often they fall into a protected shield.
Table 2 shows that 7.6% of the 50k images that MDTv2 generates without SPELL are within an $L_{2}$ distance of $r=60$ of their nearest neighbor on ImageNet. Figure 7 shows examples and verifies that such images are indeed nearly copies of existing images. Adding SPELL with $r=60$ reduces this rate down to 0.16%. Figure 7 shows that the images are indeed not too close to their ImageNet neighbors anymore. However, the runtime increases. This is not due to SPELL—in all previous experiments with $K=128$, SPELL does not increase the runtime, see Appendix G—but due to the approximate nearest neighbor search algorithm over the $K=1.2M$ images that we implement on CPU. Reducing the number of Voronoi cells that the nearest neighbor algorithm searches for shields allows to speed up the generation time, at the cost of a catching less shields. Further improvements in $L_{2}$ based neighbor search techniques will further increase the protection rate and compute overhead.
We take one final look at precision and recall in this special use-case, where the goal is to generate images that are similar but not equal to the train images. Expectedly, the recall over the validation images decreases when repelling from all training images, because validation images may fall into the shield radii around train images. However, the precision remains largely unaffected. This demonstrates again that SPELL stays on the image manifold, even when repelling from many images. Finally, the last two images in Figure 7 give more insight into the workings of the $L_{2}$ similarity in the VAE latent space that MDTv2 diffuses in. Apparently, image distances inside the VAE space encode a visual similarity where images with similar colors and compositions are close to one another. SPELL could also create shields in semantic spaces, e.g., by comparing the DINOv2 embeddings of expected image outputs, which we leave for future works.
EDMv2samplesRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionImageNetneighborRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionEDMv2 + SPELLRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionImageNetneighborRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption